# MDE-AgriVLN: Agricultural Vision-and-Language Navigation with Monocular Depth Estimation

### v1

@@ Abstract
Agricultural robots are serving as powerful assistants across a wide range of agricultural tasks, nevertheless, still heavily relying on manual operations or railway systems for movement. The AgriVLN method and the A2A benchmark pioneeringly extend Vision-and-Language Navigation (VLN) to the agricultural domain, enabling a robot to navigate to a target position following a natural language instruction. Unlike human binocular vision, most agricultural robots are only given a single camera for monocular vision, which results in limited spatial perception. To bridge this gap, we present the method of Agricultural Vision-and-Language Navigation with Monocular Depth Estimation (MDE-AgriVLN), in which we propose the MDE module generating depth features from RGB images, to assist the decision-maker on reasoning. When evaluated on the A2A benchmark, our MDE-AgriVLN method successfully increases Success Rate from 0.23 to 0.32 and decreases Navigation Error from 4.43m to 4.08m, demonstrating the state-of-the-art performance in the agricultural VLN domain.
Code: https://github.com/AlexTraveling/MDE-AgriVLN.
@@ I Introduction
To bridge this gap, we present the method of Agricultural Vision-and-Language Navigation with Monocular Depth Estimation (MDE-AgriVLN). First, we propose the MDE module, in which we employ a monocular depth estimator as the backbone and establish two representation paradigms. Second, we integrate the MDE module into the base model to establish our MDE-AgriVLN method. Third, we evaluate it on the A2A [arXiv:AgriVLN] benchmark, successfully increasing SR from 0.23 to 0.32 and decreasing NE from 4.43m to 4.08m. On the example in Fig. LABEL:fig:teaser, MDE-AgriVLN utilizes monocular depth estimation to assist its reasoning, thereby properly thinking that the robot is still on the path and approaching the blue boxes with plants, then predicting the FORWARD action, which sufficiently surpasses the performance of Baseline and narrow the gap with Human.
@@ II-C Monocular Depth Estimation
Figure 2: MDE-AgriVLN methodology illustration: The MDE module (yellow part) takes a single frame from a camera video streaming as input, to output the depth feature in two representation classifications. The base model (green part) simultaneously understand the instruction, RGB input and depth input, to reason the most appropriate low-level action with an explicit thought.
@@ III-B1 Depth Matrix
A complete depth matrix is massive and tedious. Therefore, if we directly inputed a complete depth matrix to the Vision-Language Model (VLM) in the base model, the context length would exceed the limitation. To address this issue, we innovate a simple yet effective downsampling function, denoted as $\mathcal{S}\,(\,\cdot\,)$. We denote the downsampling ratio as $r$, then the downsampled depth matrix $D_{t}^{matrix}$ is in the size of $\mathbb{H}/r\times\mathbb{W}/r$, denoted as $\mathbb{H}^{\prime}\times\mathbb{W}^{\prime}$. We uniformly split a complete frame into $\mathbb{H}^{\prime}\times\mathbb{W}^{\prime}$ dispersed patches, then only retain the central single pixel in each patch. The complete downsampling process is defined as:
@@ IV-E Qualitative Experiment
Towards a more comprehensive discussion, we share both a successful example and a failed example, as illustrated in Fig. 4. We mark all the VLM’s pivotal reasoning thoughts with italics in the text and underline in the figure.
@@ Acknowledgment
Thanks to Kota Kinabalu, Brunei, Penang and Siem Reap for the impressive traveling experiences, giving us a chilled vibe for writing. Thanks to Jingfeng Wang for being the participant offering the Human reasoning thoughts in Fig. LABEL:fig:teaser. Thanks to Yuanquan Xu, the inspiration to us.

### v2

@@ Abstract
Agricultural robots are serving as powerful assistants across a wide range of agricultural tasks, nevertheless, still heavily relying on manual operations or railway systems for movement. The AgriVLN method and the A2A benchmark pioneeringly extended Vision-and-Language Navigation (VLN) to the agricultural domain, enabling a robot to navigate to a target position following a natural language instruction. Unlike human binocular vision, most agricultural robots are only given a single camera for monocular vision, which results in limited spatial perception. To bridge this gap, we present the method of Agricultural Vision-and-Language Navigation with Monocular Depth Estimation (MDE-AgriVLN), in which we propose the MDE module generating depth features from RGB images, to assist the decision-maker on multimodal reasoning. When evaluated on the A2A benchmark, our MDE-AgriVLN method successfully increases Success Rate from 0.23 to 0.32 and decreases Navigation Error from 4.43m to 4.08m, demonstrating the state-of-the-art performance in the agricultural VLN domain. Code: https://github.com/AlexTraveling/MDE-AgriVLN.
@@ I Introduction
To bridge this gap, we present the method of Agricultural Vision-and-Language Navigation with Monocular Depth Estimation (MDE-AgriVLN). First, we propose the MDE module, in which we employ a monocular depth estimator as the backbone and establish two representation paradigms. Second, we integrate the MDE module into the base model to establish our MDE-AgriVLN method. Third, we evaluate it on the A2A [arXiv:AgriVLN] benchmark, successfully increasing SR from 0.23 to 0.32 and decreasing NE from 4.43m to 4.08m. On the example in Fig. LABEL:fig:teaser, MDE-AgriVLN utilizes monocular depth estimation to assist its reasoning, thereby properly thinking that the robot is still on the path and approaching the blue boxes with plants, then predicting the FORWARD action, which sufficiently surpasses the performance of Baseline and narrow the performance gap with Human.
@@ II-C Monocular Depth Estimation
The complete version of Related Works with more details is available in Appendix.
Figure 2: MDE-AgriVLN methodology illustration: The MDE module (yellow part) takes a single frame from a camera video streaming as input, to output the depth feature in two classifications of representation. The base model (green part) simultaneously understand the instruction, RGB input and depth input, to reason the most appropriate low-level action with an explicit thought.
@@ III-B1 Depth Matrix
A complete depth matrix is massive and tedious. Therefore, if we directly inputted a complete depth matrix to the Vision-Language Model (VLM) in the base model, the context length would exceed the limitation. To address this issue, we innovate a simple yet effective downsampling function, denoted as $\mathcal{S}\,(\,\cdot\,)$. We denote the downsampling ratio as $r$, then the downsampled depth matrix $D_{t}^{matrix}$ is in the size of $\mathbb{H}/r\times\mathbb{W}/r$, denoted as $\mathbb{H}^{\prime}\times\mathbb{W}^{\prime}$. We uniformly split a complete frame into $\mathbb{H}^{\prime}\times\mathbb{W}^{\prime}$ dispersed patches, then only retain the central single pixel in each patch. The complete downsampling process is defined as:
@@ IV-E Qualitative Experiment
Towards a more comprehensive discussion, we share both a successful example and a failed example, as illustrated in Fig. 4. We mark all the VLM’s pivotal reasoning thoughts with italics in the text and underline in the figure. The demonstration video is available in Appendix.
@@ Acknowledgment
Thanks to Kota Kinabalu, Brunei, Penang and Siem Reap for the impressive traveling experiences, giving us a chilled vibe for writing. Thanks to Kathy Wang for being the participant offering the Human reasoning thoughts in Fig. LABEL:fig:teaser. Thanks to Yuanquan Xu, the inspiration to us.